OpenStudy (anonymous):

what is the derivative of f(t)=(t^2)(e^t+1) I am very confused

OpenStudy (anonymous):

well to solve that you need to use the product rule: if y = u(t)*v(t) then y' = u'v + uv' then pose u(t) = t^2 and v(t) = (e^t+1) and follow the formula